Oriole Appliance uses a perpetual inventory system. For its flat-screen television sets, the January 1 inventory was 3 sets at $390 each. On January 10, Oriole purchased 6 units at $460 each. The company sold 2 units on January 8 and 5 units on January 15.

Compute the ending inventory under FIFO, LIFO and  ending inventory under moving-average cost
Answer #1
Unit Unit Cost Total Cost
Beginning inventory 3 390 1170
Sale Jan 8 2
Purchase Jan 10 6 460 2760
Sale Jan 15 5

Total units available for sale = 3+6 = 9 Units

Sale unit = 2+5 = 7 Units

a) FIFO = 2*460 = 920

b) LIFO = 390+460 = 850

c) Weighted average = (1*390+2760)/7*2 = 900